HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

RECENT PROGRESS IN THE FULL-DIMENSIONAL CALCULATIONS OF INTRA- AND INTERMOLECULAR ROVIBRATIONAL STATES OF

H_{2}O-HF AND H_{2}O-HCl FROM POTENTIAL ENERGY SURFACES

\underline{DOMINIKA VIGLASKA} , Groupe de Spectrométrie Moléculaire et Atmosphérique UMR CNRS 7331, UFR Sciences BP 1039, 51687 Reims Cedex 2,

France; XIAO-GANG WANG, TUCKER CARRINGTON, Chemistry Department, Queen’s University Chemistry Department, Queen’s University,

Kingston, Ontario K7L 3N6, Canada

Molecular complexes play an important role in the chemistry of planetary atmospheres. H_{2}O-HF and H_{2}O-HCl are of considerable interest because of two C_{s} equivalent minima separated by a very low inversion barrier (≈ 70 and ≈ 50 cm^{−1}, respectively). The recent methodological advances based on the full-dimensional calculations of intra - and intermolecular rovibrational states have significantly improved the understanding of molecular dynamics of Van der Waals complexes ^{1}^{,}^{2}. The rigid-monomer approximation is a good compromise when intramolecular frequencies are much higher than intermolecular frequencies but their application is often not sufficient for the high-resolution molecular spectroscopy. In particular, we show that for H_{2}O - HF dimer, the coupling between the intra- and intermolecular modes is not negligible (few cm^{−1} of difference between 5D and 9D calculations) and should be taken into account to predict accurately all the transition frequencies and tunneling splittings. All our calculations are based on a full dimensional ab initio potential energy surfaces^{3}^{,}^{4}. The kinetic energy operator discussed initially by Brocks et al.^{5} and written in terms of angular momentum operators defined in monomer- and dimer-fixed frames is used. A G_{4} symmetry-adapted Lanczos algorithm and an uncoupled product basis are employed^{1}. All computed results were confronted with available experimental data. The agreement between theory and experiment is clearly improved when 9D calculations are considered and confirms our suggestion of a possible reassignment for two bands of H_{2}O-HF. H_{2}O-HX dimers are excellent candidates for more experimental high-resolution studies.

1XG. Wang, T. Carrington, J. Chem. Phys., 2023, 158(8), 084107

2P.M. Felker, Z. Bacic, Phys.Chem. Chem. Phys., 2022, 24, 24655

3D. Viglaska, XG. Wang, T. Carrington, D.P. Tew, J.Mol. Spectrosc. 384 (2022), 111587

4Y.liu, J. Li, P.M. Felker, Z. Bacic, Phys. Chem. Chem. Phys., 2021, 23,7101

5G. Brocks et al., Mol. Phys. 50,1023 (1983)

p-number: p194  Submitted on Thu, 15 Jun 2023 16:41:19 +0200